OpenStudy (anonymous):

integrate sec^2 x/4+tan x

OpenStudy (amistre64):

the top is the derivative of the bottom

OpenStudy (amistre64):

\[\int \frac{du}{u}\]

OpenStudy (anonymous):

could you show me the steps? because i cant understand it

OpenStudy (amistre64):

im not sure what you mean by steps. this is a basic antiderivative.

OpenStudy (amistre64):

its of the from:\[\int \frac{1}{u}du\] what would that integrate to?

OpenStudy (anonymous):

ln u ?

OpenStudy (amistre64):

yes; except u = 4+tan x

OpenStudy (amistre64):

oh, and +c on the end :)

OpenStudy (amistre64):

to dbl chk; take the derivative of ln(4+tan x) and see if you get that result

OpenStudy (anonymous):

and how about this question ? integrate x^2+1/x^3+3x ?

OpenStudy (amistre64):

same concept; notice that the top is almost the derivative of the bottom

OpenStudy (amistre64):

what would ln(x^3+3x) derive to? and youll see whats missing

OpenStudy (anonymous):

do i write integral sign 1/u x du/3 ?

OpenStudy (amistre64):

that would work too ;)

OpenStudy (anonymous):

it would be then 1/3 x du/u right ?

OpenStudy (amistre64):

exactly

OpenStudy (anonymous):

then i take the 1/3 out and it would be 1/u ?

OpenStudy (amistre64):

youve got it, your like a pro by now

OpenStudy (anonymous):

then i substitute it with x^3+3x ?

OpenStudy (anonymous):

thanks ! you have helped me understand it better ! ;)

OpenStudy (amistre64):

\[\frac{1}{3}\int \frac{1}{u}du\to\ \frac{1}{3}ln(u);\ u=x^3+3x\]